# Instruction Tuning for Large Language Models: A Survey

### v7

@@ Abstract
This paper surveys research works in the quickly advancing field of instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT) [^1] In this paper, unless specified otherwise, instruction tuning (IT) will be equivalent to supervised fine-tuning (SFT)., a crucial technique to enhance the capabilities and controllability of large language models (LLMs). Instruction tuning refers to the process of further training LLMs on a dataset consisting of (instruction, output) pairs in a supervised fashion, which bridges the gap between the next-word prediction objective of LLMs and the users’ objective of having LLMs adhere to human instructions. In this work, we make a systematic review of the literature, including the general methodology of IT, the construction of IT datasets, the training of IT models, and applications to different modalities, domains and application, along with analysis on aspects that influence the outcome of IT (e.g., generation of instruction outputs, size of the instruction dataset, etc). We also review the potential pitfalls of IT along with criticism against it, along with efforts pointing out current deficiencies of existing strategies and suggest some avenues for fruitful research. footnotetext: Zhejiang University, Shannon.AI, Nanyang Technological University, Amazon footnotetext: Email: sy_zhang@zju.edu.cn footnotetext: Project page can be found at: https://github.com/xiaoya-li/Instruction-Tuning-Survey footnotetext: * The latest update was on October 16, 2024.
@@ 1 Introduction
The field of large language models (LLMs) has witnessed remarkable progress in recent years. LLMs such as GPT-3 (Brown et al., 2020b), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023a) have demonstrated impressive capabilities across a wide range of natural language tasks (Zhao et al., 2021; Wang et al., 2022b, 2023b; Wan et al., 2023; Sun et al., 2023c; Wei et al., 2023a; Li et al., 2023a; Gao et al., 2023a; Yao et al., 2023; Yang et al., 2022a; Qian et al., 2022; Lee et al., 2022; Yang et al., 2022b; Gao et al., 2023b; Ning et al., 2023; Liu et al., 2021b; Wiegreffe et al., 2021; Sun et al., 2023b, a; Adlakha et al., 2023; Chen et al., 2023b). One of the major issues with LLMs is the mismatch between the training objective and users’ objective: LLMs are typically trained on minimizing the contextual word prediction error on large corpora; while users want the model to "follow their instructions helpfully and safely" (Radford et al., 2019; Brown et al., 2020a; Fedus et al., 2021; Rae et al., 2021; Thoppilan et al., 2022)
@@
To address this mismatch, instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT), is proposed, serving as an effective technique to enhance the capabilities and controllability of large language models. It involves further training LLMs using (instruction, output) pairs, where instruction denotes the human instruction for the model, and output denotes the desired output that follows the instruction. The benefits of IT are threefold: (1) Finetuning an LLM on the instruction dataset bridges the gap between the next-word prediction objective of LLMs and the users’ objective of instruction following; (2) IT allows for a more controllable and predictable model behavior compared to standard LLMs. The instructions serve to constrain the model’s outputs to align with the desired response characteristics or domain knowledge, providing a channel for humans to intervene with the model’s behaviors; and (3) IT is computationally efficient and can help LLMs rapidly adapt to a specific domain without extensive retraining or architectural changes.
@@
Despite its effectiveness, IT also poses challenges: (1) Crafting high-quality instructions that properly cover the desired target behaviors is non-trivial: existing instruction datasets are usually limited in quantity, diversity, and creativity; (2) there has been an increasing concern that IT only improves on tasks that are heavily supported in the IT training dataset Gudibande et al. (2023); and (3) there has been an intense criticism that IT only captures surface-level patterns and styles (e.g., the output format) rather than comprehending and learning the task Kung and Peng (2023). Improving instruction adherence and handling unanticipated model responses remain open research problems. These challenges highlight the importance of further investigations, analysis, and summarization in this field, to optimize the fine-tuning process and better understand the behavior of instruction tuned LLMs.
@@
  • Section 3 outlines the construction process of commonly-used IT representative datasets.
@@
  • Section 6 reviews efforts to adapt LLMs to different domains and applications using the IT strategy.
@@
  • Section 8 presents the evaluation of IT models, analysis on them, along with criticism against them.
@@ 2.1 Instruction Dataset Construction
For multi-turn conversational IT datasets, we can have large language models self-play different roles (user and AI assistant) to generate messages in a conversational format  (Xu et al., 2023b).
@@ 2.2 Instruction Tuning / Supervised Fine-tuning
Based on the collected IT dataset, a pretrained model can be directly fune-tuned in a fully-supervised manner, where given the instruction and the input, the model is trained by predicting each token in the output sequentially.
@@ 3 Datasets
In this section, we detail instruction tuning datasets in the community, categorizing them into three classes: (1) Human-crafted Data, (2) Synthetic Data via Distillation, and (3) Synthetic Data via Self-improvement. Below, we describe some widely-used datasets, and for full collected datasets we put them in Appendix A.
@@ 3.2 Synthetic Data via Distillation
[[[[Human]]]] Hello!
[[[[AI]]]] Hi! How can I help you?
@@ 4.11 Others
(Wang et al., 2023d) is a large language model trained by fine-tuning OPT (6.7B) (Zhang et al., 2022a) on a mixed instruction dataset, which contains FLAN V2 (Longpre et al., 2023), CoT (Wei et al., 2022), Dolly (Conover et al., 2023a), Open Assistant-1131313https://huggingface.co/datasets/OpenAssistant/oasst1, GPT4-Alpaca [^14] https://huggingface.co/datasets/vicgalle/alpaca-gpt4, Code-Alpaca (Chaudhary, 2023), and ShareGPT [^15] https://sharegpt.com/. After fine-tuning, TÜLU (6.7B) reaches on average 83% of ChatGPT’s performance and 68% of GPT-4’s performance.
@@ 5.1 Multi-modality Datasets
(Zhang et al., 2023c) is a large-scale medical visual question-answering dataset that comprises 227k image-question pairs of 149k images, covering various modalities or diseases. The dataset can be used for both open-ended and multiple-choice tasks. The pipeline for generating the PMC-VQA dataset involves collecting image-caption pairs from the PMC-OA (Lin et al., 2023b) dataset, using ChatGPT to generate question-answer pairs, and manually verifying a subset of the dataset for quality. The authors propose a generative-based model MedVInT for medical visual understanding by aligning visual information with a large language model. MedVInT pretrained on PMC-VQA achieves state-of-the-art performance and outperforms existing models on VQA-RAD (Lau et al., 2018) and SLAKE (Liu et al., 2021a) benchmarks, with 81.6% accuracy on VQA-RAD and 88.0% accuracy on SLAKE.
@@
(Xu et al., 2024) is the largest public-available human-annotated visual instruction tuning dataset that consists of 1,664,261 instances and 200+ diverse vision-language tasks derived from 101 open-source computer vision datasets. Each task is accompanied by expertly written instructions and meticulously crafted templates for inputs and outputs. The dataset covers a broad spectrum of tasks, including image captioning, visual question-answering, and visual comprehension. Designed to enhance research and application in vision-language model domains, Vision-Flan aims to expand the horizons of interaction and comprehension between visual and linguistic modalities. It provides researchers and developers with a valuable resource to push the envelope of vision-language models and to innovate algorithms across a diverse array of fields.
@@ 6.2 Intent Classification and Slot Tagging
Figure 10: The overview framework of InstructUIE. The figure is copied from Wang et al. (2023c).
@@ 6.3 Information Extraction
(Wang et al., 2023c) is a unified information extraction (IE) framework based on instruction tuning, which transforms IE tasks to the seq2seq format and solves them by fine-tuning 11B FlanT5 (Chung et al., 2022) on the constructed IT dataset. Figure 10 shows the overall architecture of InstructUIE. It introduces IE INSTRUCTIONS, a benchmark of 32 diverse information extraction datasets in a unified text-to-text format with expert-written instructions. Each task instance is delineated by four properties: task instruction, options, text, and output. Task instruction contains information such as the type of information to be extracted, the output structure format, and additional constraints or rules that need to be adhered to during the extraction process. Options refer to the output label constraints of a task. Text refers to the input sentence. Output is the sentence obtained by converting the original tags of the sample (e.g. "entity tag: entity span" for NER). In the supervised setting, InstructUIE performs comparably to BERT (Devlin et al., 2018) and outperforms the state-of-the-art and GPT3.5 (Brown et al., 2020a) in zero-shot settings.
@@ 8.1 Close-ended Evaluations
8.2 GPT-Based Evaluations
GPT-based evaluations refer to methods that leverage GPT-4 OpenAI (2023) to assess the output of LLMs. There are three primary reasons for this approach: (1) Efficiency – Manually reviewing numerous LLM outputs can be labor-intensive, whereas GPT-4 can evaluate large-scale responses quickly, saving both time and effort; (2) Reliable Benchmark – As one of the most advanced models available, GPT-4 provides a dependable benchmark, allowing researchers to compare the performance of different LLMs against a high standard; and (3) Enhanced Capability – With improved comprehension and reasoning over previous models, GPT-4 is better suited to analyze subtle aspects of language generation and handle complex outputs from other LLMs. In the following, we detail 4 commonly accepted GPT-based evaluations:
@@ 8.2 GPT-Based Evaluations
8.3 HELM Evaluation
(1) Broad coverage.
(2) Multi-metric measurement.
(3) Standardization.
@@ 8.4 Low-resource Instruction Tuning
Gupta et al. (2023) attempts to estimate the minimal downstream training data required by IT models to match the SOTA supervised models over various tasks. Gupta et al. (2023) conducted experiments on 119 tasks from Super Natural Instructions (SuperNI) in both single-task learning (STL) and multi-task learning (MTL) settings. The results indicate that in the STL setting, IT models with only 25% of downstream training data outperform the SOTA models on those tasks, while in the MTL setting, just 6% of downstream training data can lead IT models to achieve the SOTA performance. These findings suggest that instruction tuning can effectively assist a model in quickly learning a task even with limited data.
However, due to resource limitations, Gupta et al. (2023) did not conduct experiments on LLMs, like T5-11B. So, to gain a more comprehensive understanding of the IT models, further investigation using larger language models and datasets is necessary.

### v9

@@ Abstract
This paper surveys research works in the quickly advancing field of instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT) [^1] In this paper, unless specified otherwise, supervised fine-tuning (SFT) and instruction tuning (IT) are used interchangeably. , a crucial technique to enhance the capabilities and controllability of large language models (LLMs). Instruction tuning refers to the process of further training LLMs on a dataset consisting of (instruction, output) pairs in a supervised fashion, which bridges the gap between the next-word prediction objective of LLMs and the users’ objective of having LLMs adhere to human instructions. In this work, we make a systematic review of the literature, including the general methodology of SFT, the construction of SFT datasets, the training of SFT models, and applications to different modalities, domains and application, along with analysis on aspects that influence the outcome of SFT (e.g., generation of instruction outputs, size of the instruction dataset, etc). We also review the potential pitfalls of SFT along with criticism against it, along with efforts pointing out current deficiencies of existing strategies and suggest some avenues for fruitful research. footnotetext: Zhejiang University, Shannon.AI, Nanyang Technological University, Amazon footnotetext: Email: sy_zhang@zju.edu.cn footnotetext: Project page can be found at: https://github.com/xiaoya-li/Instruction-Tuning-Survey footnotetext: * The latest update was on Aug. 11, 2025 (Version 6).
@@ 1 Introduction
The field of large language models (LLMs) has witnessed remarkable progress in recent years. LLMs such as GPT-3 (Brown et al., 2020b), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023a) have demonstrated impressive capabilities across a wide range of natural language tasks (Zhao et al., 2021; Wang et al., 2022b, 2023c; Wan et al., 2023; Sun et al., 2023c; Wei et al., 2023a; Li et al., 2023a; Gao et al., 2023a; Yao et al., 2023; Yang et al., 2022a; Qian et al., 2022; Lee et al., 2022; Yang et al., 2022b; Gao et al., 2023b; Ning et al., 2023; Liu et al., 2021b; Wiegreffe et al., 2021; Sun et al., 2023b, a; Adlakha et al., 2023; Chen et al., 2023b). One of the major issues with LLMs is the mismatch between the training objective and users’ objective: LLMs are typically trained on minimizing the contextual word prediction error on large corpora; while users want the model to "follow their instructions helpfully and safely" (Radford et al., 2019; Brown et al., 2020a; Fedus et al., 2021; Rae et al., 2021; Thoppilan et al., 2022)
@@
To address this mismatch, instruction tuning (IT), which can also be referred to as supervised fine-tuning (SFT), is proposed, serving as an effective technique to enhance the capabilities and controllability of large language models. It involves further training LLMs using (instruction, output) pairs, where instruction denotes the human instruction for the model, and output denotes the desired output that follows the instruction. The benefits of SFT are threefold: (1) Finetuning an LLM on the instruction dataset bridges the gap between the next-word prediction objective of LLMs and the users’ objective of instruction following; (2) SFT allows for a more controllable and predictable model behavior compared to standard LLMs. The instructions serve to constrain the model’s outputs to align with the desired response characteristics or domain knowledge, providing a channel for humans to intervene with the model’s behaviors; and (3) SFT is computationally efficient and can help LLMs rapidly adapt to a specific domain without extensive retraining or architectural changes.
@@
Despite its effectiveness, SFT also poses challenges: (1) Crafting high-quality instructions that properly cover the desired target behaviors is non-trivial: existing instruction datasets are usually limited in quantity, diversity, and creativity; (2) there has been an increasing concern that SFT only improves on tasks that are heavily supported in the SFT training dataset Gudibande et al. (2023); and (3) there has been an intense criticism that SFT only captures surface-level patterns and styles (e.g., the output format) rather than comprehending and learning the task Kung and Peng (2023). Improving instruction adherence and handling unanticipated model responses remain open research problems. These challenges highlight the importance of further investigations, analysis, and summarization in this field, to optimize the fine-tuning process and better understand the behavior of instruction tuned LLMs.
@@
  • Section 3 outlines the construction process of commonly-used SFT representative datasets, along with multi-step reasoning datasets designed to enhance LLM performance on complex reasoning tasks such as mathematics and coding.
@@
  • Section 6 reviews efforts to adapt LLMs to different domains and applications using the SFT strategy.
@@
  • Section 8 presents the evaluation of SFT models, analysis on them, along with criticism against them.
  • Section 9 analyzes the role of SFT in comparison with recent, highly effective reinforcement learning–based methods (e.g., RLHF, DPO, and GRPO).
@@ 2.1 Instruction Dataset Construction
For multi-turn conversational SFT datasets, we can have large language models self-play different roles (user and AI assistant) to generate messages in a conversational format  (Xu et al., 2023b).
@@ 2.2 Instruction Tuning / Supervised Fine-tuning
Based on the collected SFT dataset, a pretrained model can be directly fune-tuned in a fully-supervised manner, where given the instruction and the input, the model is trained by predicting each token in the output sequentially.
@@ 3 Datasets
In this section, we detail instruction tuning datasets in the community, categorizing them into three classes: (1) Human-crafted Data, (2) Synthetic Data via Distillation, and (3) Synthetic Data via Self-improvement. Further more, in light of the impressive performance of recent multi-step reasoning LLMs (e.g., OpenAI o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025)), this section also presents a detailed overview of how reasoning datasets are constructed. These datasets, typically built using one or a combination of the three strategies mentioned above, are specifically designed to enhance LLMs’ multi-step thinking capabilities. Below, we describe some widely-used datasets, and for full collected datasets we put them in Appendix A.
@@ 3.2 Synthetic Data via Distillation
[[Human]] Hello!
[[AI]] Hi! How can I help you?
@@ 3.3.2 Instruction Back-translation
3.4 Reasoning Datasets
Reasoning datasets focus on logical progression, multi-step thinking, and structured problem-solving. By incorporating challenging problems, well-defined scenarios, and diverse contexts, they help bridge the gap between generic text data, that most LLMs are trained on, and specialized reasoning skills. In this section, we briefly review several reasoning-formatted datasets, with the full list provided in Appendix A.
3.4.1 PRM800K
PRM800K Lightman et al. (2023) is a large-scale, open-source dataset containing step-level human feedback labels, created through a combination of machine-generated and human-generated methods. It comprises 800K annotated steps from 75K solutions to 12K problems sourced from the MATH Hendrycks et al. (2021) dataset. Each entry includes two components: (1) steps—intermediate reasoning steps generated sequentially by GPT-4, and (2) labels—human annotations marking each step as correct (positive), incorrect (negative), or ambiguous (neutral). The dataset was built through three stages: (1) GPT-4 generated step-by-step solutions to MATH problems; (2) only solutions with correct final answers were retained; and (3) human annotators labeled each step, with special attention to ’convincing wrong-answer’ cases, high-quality but incorrect solutions (Figure 7), to maximize feedback value.
Figure 7: A screenshot of the interface used to collect feedback in PRM800KLightman et al. (2023). The figure is borrowed from  Lightman et al. (2023).
3.4.2 O1-Journey
O1-Journey Qin et al. (2024) is an open-source English reasoning dataset with 677 instances, 327 of which are used for training. Built through a mix of machine- and human-generated methods, each instance includes a question (the problem to solve), an answer (the correct solution), and a longCOT, a detailed chain-of-thought incorporating intermediate steps, reflections, and corrections. Its construction involves three stages: (1) Reasoning Tree Generation, a pre-trained policy model produces reasoning trees for problems from MATHHendrycks et al. (2021) and PRM800KLightman et al. (2023), which are then evaluated by a reward model, with incorrect trees discarded; (2) Reasoning Data Expansion, a multi-agent system generates reasoning steps, with one agent producing solutions and another providing feedback in an iterative process to emulate human-like reflection and revision; and (3) Data Augmentation, human annotators manually refine and enhance the expanded reasoning data.
3.4.3 MathGenie
MathGenie Lu et al. (2024) is a dataset created to produce synthetic math problems that enhance large language models’ mathematical reasoning capabilities. The resulting corpus, MathGenieData, contains 170K question–solution pairs, 110K from GSM8K Cobbe et al. (2021) and 60K from MATH Hendrycks et al. (2021), and is used for fine-tuning various pre-trained models. Its construction follows a three-stage pipeline (Figure 8): (1) Iterative Solution Augmentation, starting with a 15K problem seed set from GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), a fine-tuned LLaMA-2 70B Touvron et al. (2023b) model generates diverse alternative solutions that depart significantly from the originals; (2) Question Back-Translation, the fine-tuned LLaMA-2 70B Touvron et al. (2023b) converts these augmented solutions into new math questions, guided by solution constraints to ensure validity and relevance; and (3) Verification-Based Filtering, a code-integrated solution generator, also fine-tuned on LLaMA-2 70B Touvron et al. (2023b), produces solutions for the new questions, which are rigorously verified through combined natural-language and code reasoning to retain only correct results.
Figure 8: Framework of MathGenie Lu et al. (2024). Step 1: The Iterative Solution Augmentation method adds more examples to human-annotated solutions in the GSM8K and MATH datasets. Step 2: Question Back-translation turns these solutions into new questions. Step 3: Verification-Based Solution Filtering selects reliable code-based solutions by generating and verifying them through a series of validation steps. The figure is borrowed from  Lu et al. (2024).
3.4.4 DeepSeekMath
The DeepSeekMath Corpus Shao et al. (2024) is a large-scale, open-source dataset for mathematical reasoning, comprising 120 billion tokens generated through both machine and human efforts. Its core source is “Common Crawl”, supplemented with material from “AlgebraicStack”, “arXiv”, “GitHub”, and other natural language texts. Aimed at improving language models’ mathematical reasoning abilities, it is multilingual, with a strong emphasis on English and Chinese math content.
Construction follows an iterative collect–refine cycle (Figure 9): (1) Classifier Training, a fastText-based model is first trained with OpenWebMath Paster et al. (2023) as positive examples and a range of general web pages as negatives, with retraining after each new round of data collection; (2) Mathematics Extraction, the classifier identifies additional math-rich content from “Common Crawl”, which is then refined through human annotation. To ensure quality and prevent benchmark contamination, pages containing known benchmark Q&A are removed. This loop steadily improves classifier precision while expanding the dataset’s scope
Figure 9: Pipeline of DeepSeekMath Shao et al. (2024). The iterative process for gathering math-related web pages from Common Crawl. The figure is borrowed from  Shao et al. (2024).
@@ 4.11 Others
(Wang et al., 2023e) is a large language model trained by fine-tuning OPT (6.7B) (Zhang et al., 2022a) on a mixed instruction dataset, which contains FLAN V2 (Longpre et al., 2023), CoT (Wei et al., 2022), Dolly (Conover et al., 2023a), Open Assistant-1131313https://huggingface.co/datasets/OpenAssistant/oasst1, GPT4-Alpaca [^14] https://huggingface.co/datasets/vicgalle/alpaca-gpt4, Code-Alpaca (Chaudhary, 2023), and ShareGPT [^15] https://sharegpt.com/. After fine-tuning, TÜLU (6.7B) reaches on average 83% of ChatGPT’s performance and 68% of GPT-4’s performance.
@@ 5.1 Multi-modality Datasets
(Zhang et al., 2023c) is a large-scale medical visual question-answering dataset that comprises 227k image-question pairs of 149k images, covering various modalities or diseases. The dataset can be used for both open-ended and multiple-choice tasks. The pipeline for generating the PMC-VQA dataset involves collecting image-caption pairs from the PMC-OA (Lin et al., 2023c) dataset, using ChatGPT to generate question-answer pairs, and manually verifying a subset of the dataset for quality. The authors propose a generative-based model MedVInT for medical visual understanding by aligning visual information with a large language model. MedVInT pretrained on PMC-VQA achieves state-of-the-art performance and outperforms existing models on VQA-RAD (Lau et al., 2018) and SLAKE (Liu et al., 2021a) benchmarks, with 81.6% accuracy on VQA-RAD and 88.0% accuracy on SLAKE.
@@
(Xu et al., 2024b) is the largest public-available human-annotated visual instruction tuning dataset that consists of 1,664,261 instances and 200+ diverse vision-language tasks derived from 101 open-source computer vision datasets. Each task is accompanied by expertly written instructions and meticulously crafted templates for inputs and outputs. The dataset covers a broad spectrum of tasks, including image captioning, visual question-answering, and visual comprehension. Designed to enhance research and application in vision-language model domains, Vision-Flan aims to expand the horizons of interaction and comprehension between visual and linguistic modalities. It provides researchers and developers with a valuable resource to push the envelope of vision-language models and to innovate algorithms across a diverse array of fields.
@@ 6.2 Intent Classification and Slot Tagging
Figure 13: The overview framework of InstructUIE. The figure is copied from Wang et al. (2023d).
@@ 6.3 Information Extraction
(Wang et al., 2023d) is a unified information extraction (IE) framework based on instruction tuning, which transforms IE tasks to the seq2seq format and solves them by fine-tuning 11B FlanT5 (Chung et al., 2022) on the constructed SFT dataset. Figure 13 shows the overall architecture of InstructUIE. It introduces IE INSTRUCTIONS, a benchmark of 32 diverse information extraction datasets in a unified text-to-text format with expert-written instructions. Each task instance is delineated by four properties: task instruction, options, text, and output. Task instruction contains information such as the type of information to be extracted, the output structure format, and additional constraints or rules that need to be adhered to during the extraction process. Options refer to the output label constraints of a task. Text refers to the input sentence. Output is the sentence obtained by converting the original tags of the sample (e.g. "entity tag: entity span" for NER). In the supervised setting, InstructUIE performs comparably to BERT (Devlin et al., 2018) and outperforms the state-of-the-art and GPT3.5 (Brown et al., 2020a) in zero-shot settings.
@@ 8.1 Close-ended Evaluations
8.2 HELM Evaluation
(1) Broad coverage.
(2) Multi-metric measurement.
(3) Standardization.
8.3 LLM As a Judge
LLM as a judge refers to a set of methods that utilize powerful LLMs, particularly GPT-4 OpenAI (2023), to evaluate the outputs of other LLMs. There are three primary reasons for this approach: (1) Efficiency – Manually reviewing numerous LLM outputs can be labor-intensive, whereas GPT-4 can evaluate large-scale responses quickly, saving both time and effort; (2) Reliable Benchmark – As one of the most advanced models available, GPT-4 provides a dependable benchmark, allowing researchers to compare the performance of different LLMs against a high standard; and (3) Enhanced Capability – With improved comprehension and reasoning over previous models, GPT-4 is better suited to analyze subtle aspects of language generation and handle complex outputs from other LLMs. In the following, we detail 4 commonly accepted judge benchmarks:
@@ 8.4 Low-resource Instruction Tuning
Gupta et al. (2023) attempts to estimate the minimal downstream training data required by SFT models to match the SOTA supervised models over various tasks. Gupta et al. (2023) conducted experiments on 119 tasks from Super Natural Instructions (SuperNI) in both single-task learning (STL) and multi-task learning (MTL) settings. The results indicate that in the STL setting, SFT models with only 25% of downstream training data outperform the SOTA models on those tasks, while in the MTL setting, just 6% of downstream training data can lead SFT models to achieve the SOTA performance. These findings suggest that instruction tuning can effectively assist a model in quickly learning a task even with limited data.
However, due to resource limitations, Gupta et al. (2023) did not conduct experiments on LLMs, like T5-11B. So, to gain a more comprehensive understanding of the SFT models, further investigation using larger language models and datasets is necessary.